# On $\mathrm{F}$-spaces of almost-Lebesgue functions

### v2

@@ 6. Approximation and separability
7. Marcinkiewicz interpolation theorems
In this section we establish an interpolation result for the spaces $\Lambda_{p}(\mathrm{X})$ in the spirit of the Marcinkiewicz interpolation theorem. This classical theorem provides strong $L_{p}$-bounds for subadditive operators assuming weaker estimates at endpoints $q$ and $r$ such that $1\leq q<p<r\leq\infty$. We will focus on the case $r=\infty$ as it is the version that can be adapted straightforwardly to our setting. An operator $T:\mathrm{dom}(T)\subseteq L_{0}(\mathrm{X})\to L_{0}(\mathrm{X})$ is called subadditive if $|T(f+g)|\leq|T(f)|+|T(g)|$ for all measurable functions $f$ and $g$ in its domain $\mathrm{dom}(T)$.
Theorem 7.1 (Marcinkiewicz interpolation theorem [11]).
Let $(\mathrm{X},\mu)$ be a $\sigma$-finite measure space and $1\leq q<\infty$. Let $T:L_{q}(\mathrm{X})+L_{\infty}(\mathrm{X})\to L_{0}(\mathrm{X})$ be a subadditive operator and assume that there exist $A,B>0$ such that
$\sup_{\lambda>0}\lambda^{q}D(T(f),\lambda)\leq A^{q}\|f\|_{q}\qquad\forall f\in L_{q}(\mathrm{X})$ (7.1)
and
$\|T(f)\|_{\infty}\leq B\,\|f\|_{\infty}\qquad\forall f\in L_{\infty}(\mathrm{X}).$ (7.2)
Then, for every $p>q$,
$\|T(f)\|_{p}\leq C\,\|f\|_{p}\qquad\forall f\in L_{p}(\mathrm{X})$ (7.3)
where
$C=2\left(\frac{p}{p-q}\right)^{\frac{1}{p}}A^{\frac{q}{p}}B^{1-\frac{q}{p}}.$ (7.4)
Here, $L_{\infty}(\mathrm{X})$ denotes the standard Lebesgue space of essentially bounded measurable functions with norm $\|\cdot\|_{\infty}$, and $D(f,\cdot):[0,\infty)\to[0,\infty]$ is the distribution function of a measurable function $f$ given by $D(f,\lambda)=\mu(|f|>\lambda)$.
In order to obtain an analogous result for the spaces $\Lambda_{p}(\mathrm{X})$, one starts by defining for any measurable function $f$ its restricted distribution function $\rho(f,\cdot):[0,1]\to[0,\infty]$ by
$\rho(f,\lambda)=D(\min(|f|,1),\lambda)=\mu\big{(}\min(|f|,1)>\lambda\big{)}.$
If $\mathrm{X}$ is $\sigma$-finite, then for each measurable function $f$ and $p\geq 1$, Tonelli’s theorem yields that
$p\int_{0}^{1}\lambda^{p-1}\rho(f,\lambda)\,\mathrm{d}\lambda=\|f\|_{\alpha_{p}}^{p}.$ (7.5)
Also, for each $\lambda\in(0,1)$ we have
$\rho(f,\lambda)\leq\frac{\|f\|_{\alpha_{p}}^{p}}{\lambda^{p}}.$ (7.6)
Lemma 7.2.
Let $f$ be a measurable function and $\lambda\in(0,1)$. Then $f$ can be decomposed as follows
$f=f\chi_{\{\min(|f|,1)>\lambda\}}+f\chi_{\{|f|\leq\lambda\}}.$ (7.7)
Proof.
Simply note that for $\lambda\in(0,1)$ it holds that
$\{\min(|f|,1)>\lambda\}=\{|f|>\lambda\}$ (7.8)
for each measurable function $f$. ∎
Identity (7.8) yields that the restricted distribution function of a measurable function $f$ equals the distribution function of $f$ restricted to the interval $(0,1)$; that is, $\rho(f,\cdot)=D(f,\cdot)|_{(0,1)}$.
Lemma 7.3.
Let $1\leq q<r<\infty$. Then, for each $p\in(q,r)$ we have
$\Lambda_{p}(\mathrm{X})\subseteq\Lambda_{q}(\mathrm{X})+\Lambda_{r}(\mathrm{X})\cap L_{\infty}(\mathrm{X}).$
Proof.
Let $f\in\Lambda_{p}(\mathrm{X})$ for some $p\in(q,r)$, and fix $\lambda>0$. We write $f=g_{\lambda}+b_{\lambda}$, where $g_{\lambda}=f\chi_{\{|f|>\lambda\}}$ and $b_{\lambda}=f\chi_{\{|f|\leq\lambda\}}$. The result follows from noting that $g_{\lambda}\in\Lambda_{q}(\mathrm{X})$ and $b_{\lambda}\in\Lambda_{r}(\mathrm{X})\cap L_{\infty}(\mathrm{X})$. ∎
The next result provides a sufficient condition for a measurable function belonging to the intersection of two almost-Lebesgue spaces to also belong to Lebesgue spaces.
Proposition 7.4.
Let $(\mathrm{X},\mu)$ be a $\sigma$-finite measure space and let $f\in\Lambda_{q}(\mathrm{X})\,\cap\,\Lambda_{r}(\mathrm{X})$ with $1\leq q<r<\infty$. Suppose that for every $\varepsilon>0$ there exists $\delta_{\varepsilon}$ such that
$\sup_{\lambda>0}\lambda\max\left(\mu(|f|\chi_{E}>\lambda)^{\frac{1}{q}},\mu(|f|\chi_{E}>\lambda)^{\frac{1}{r}}\right)<\varepsilon$ (7.9)
for each measurable set $E$ with $\mu(E)<\delta_{\varepsilon}$. Then $f\in L_{p}(\mathrm{X})$ for every $p\in(q,r)$.
Proof.
It is clear that $\Lambda_{q}(\mathrm{X})\,\cap\,\Lambda_{r}(\mathrm{X})\subseteq\Lambda_{p}(\mathrm{X})$ for every $p\in(q,r)$. Moreover, if $f$ satisfies (7.9) then $f$ is has an absolutely continuous $p$-integral for every $p\in(q,r)$. Thus, the result follows by Proposition 2.3. ∎
The previous result shows that the almost-Lebesgue spaces are not very well suited for $L_{p}$-interpolation, as one needs the extra condition (7.9) on a function in $\Lambda_{q}(\mathrm{X})\cap\Lambda_{r}(\mathrm{X})$ to deduce that it belongs to $L_{p}(\mathrm{X})$ for $p\in(q,r)$. Nevertheless, a version of Theorem 7.1 for the present setting still holds. Its proof is a simple adaptation of the proof of Theorem 7.1 in [11, Theorem 1.3.3], with minor modifications, and is included here for convenience.
Theorem 7.5.
Let $(\mathrm{X},\mu)$ be a $\sigma$-finite measure space and $1\leq q<\infty$. Let $T:\Lambda_{q}(\mathrm{X})+L_{\infty}(\mathrm{X})\to L_{0}(\mathrm{X})$ be a subadditive operator and assume that there exist $A>0$ such that
$\sup_{0<\lambda<1}\lambda^{q}\rho(T(f),\lambda)\leq A^{q}\|f\|_{\alpha_{q}}^{q}\qquad\forall f\in\Lambda_{q}(\mathrm{X}),$ (7.10)
and (7.2) holds for some $B\geq 1/2$. Then, for every $p>q$,
$\|T(f)\|_{\alpha_{p}}\leq C\,\|f\|_{\alpha_{p}}\qquad\forall f\in\Lambda_{p}(\mathrm{X})$ (7.11)
where $C$ is as in (7.4).
Proof.
Let $f\in\Lambda_{p}(\mathrm{X})$ for some $p>q$. We use the decomposition (7.7) with $\lambda_{t}=t/(2B)$, $t\in(0,1)$. The hypothesis $B\geq 1/2$ is imposed to ensure that $\lambda_{t}$ belongs to $(0,1)$ for all $t\in(0,1)$. Set
$g_{t}=f\chi_{\{\min(|f|,1)>t/(2B)\}}\quad\text{and}\quad b_{t}=f\chi_{\{|f|\leq t/(2B)\}}.$
It is clear that $\|b_{t}\|_{\infty}\leq t/(2B)$, and hence, by (7.2) we deduce that
$|T(b_{t})(x)|\leq\|T(b_{t})\|_{\infty}\leq B\|b_{t}\|_{\infty}\leq t/2$
for almost every $x\in X$, and so $\rho(T(b_{t}),t/2)=\mu(|T(b_{t})|>t/2)=0$, where we used identity (7.8). Therefore, from the subadditivity of $T$ it follows that
$\rho(T(f),t)\leq\rho(T(g_{t}),t/2)+\rho(T(b_{t}),t/2)=\rho(T(g_{t}),t/2).$
Moreover, using (7.10) with $\lambda=t/2$ leads to
$\rho(T(g_{t}),t/2)\leq 2^{q}A^{q}t^{-q}\int_{\{2B\min(|f|,1)>t\}}\min(|f|,1)^{q}\,\mathrm{d}\mu.$
Combining the previous expressions with (7.5) and using Tonelli’s theorem to interchange the order of integration yields
$\|T(f)\|_{\alpha_{p}}^{p}$ $=p\int_{0}^{1}t^{p-1}\rho(T(f),t)\,\mathrm{d}t$
$\leq p2^{q}A^{q}\int_{0}^{1}t^{p-q-1}\int_{\{2B\min(|f|,1)>t\}}\min(|f|,1)^{q}\,\mathrm{d}\mu\,\mathrm{d}t$
$=p2^{q}A^{q}\int_{X}\int_{0}^{2B\min(|f|,1)}t^{p-q-1}\min(|f|,1)^{q}\,\mathrm{d}t\,\mathrm{d}\mu$
$=\frac{p}{p-q}2^{p}A^{q}B^{p-q}\|f\|_{\alpha_{p}}^{p}$
which finishes the proof. ∎
Remark 7.6.
Note that the left-hand side of condition (7.10) is obtained from the left-hand side of condition (7.1) by restricting the values of $\lambda$ considered in the supremum, while enlarging the set of functions that satisfy it. Additionally, even if an operator satisfies condition (7.10), restricting it to functions only in $L_{q}(\mathrm{X})$ does not necessarily imply condition (7.1), as the supremum in that case is taken over a larger interval.
8. Hardy-Littlewood maximal operators
Consider $\mathrm{X}=\mathbb{R}^{d}$ with the Borel $\sigma$-algebra and the Lebesgue measure. We denote by $|E|$ the Lebesgue measure of a measurable set $E\subseteq\mathbb{R}^{d}$.
Given a measurable function $f$ on $\mathbb{R}^{d}$, recall its uncentered Hardy-Littlewood maximal function $M(f)$ given by
$M(f)(x)=\sup_{B\ni x}\frac{1}{|B|}\int_{B}|f(y)|\,\mathrm{d}y$
where the supremum is taken over all open balls that contain $x$. For each measurable function $f$ on $\mathbb{R}^{d}$ and every $\lambda>0$ it holds that [11, Theorem 1.4.6]:
$\lambda|\{M(f)>\lambda\}|\leq 3^{d}\int_{\{M(f)>\lambda\}}|f(y)|\,\mathrm{d}y.$ (8.1)
The aim of this section is to provide an application of Theorem 7.5 to establish a boundedness result for a maximal operator suited for the present setting. Given a measurable function $f$, we define its bounded maximal function $M_{b}(f)$ by
$M_{b}(f)=M(\min(|f|,1))=\sup_{B\ni x}\frac{1}{|B|}\|f\|_{\alpha}.$
Clearly, $\|M_{b}(f)\|_{\infty}\leq\min(\|f\|_{\infty},1)$ for $f\in L_{\infty}(\mathbb{R}^{d})$.
Theorem 8.1.
For every $p>1$ and $f\in\Lambda_{p}(\mathbb{R}^{d})$ we have
$\|M_{b}(f)\|_{\alpha_{p}}\leq C\,\|f\|_{\alpha_{p}}$ (8.2)
where
$C=2\left(\frac{p}{p-1}\right)^{\frac{1}{p}}3^{\frac{d}{p}}.$
Proof.
It suffices to show that the operator $M_{b}$ satisfies (7.10) with $q=1$ for some $A>0$. Let $\lambda\in(0,1)$ and note that by (7.8) we have $\{\min(M_{b}(f),1)>\lambda\}=\{M(\min(|f|,1))>\lambda\}$. Thus, for $\lambda\in(0,1)$, using (8.1) we deduce
$\lambda\rho(M_{b}(f),\lambda)$ $=\lambda|\{M(\min(|f|,1)>\lambda)\}|\leq 3^{d}\|f\|_{\alpha}.$
The result follows upon applying Theorem 7.5 with $A=3^{d}$ and $B=1$. ∎
Remark 8.2.
Alternatively, instead of using Theorem 7.5 to prove the previous result, one could apply the maximal function theorem, which states that $\|M(f)\|_{p}\leq C\|f\|_{p}$ for every $f\in L_{p}(\mathrm{X})$; a result derived using (8.1) and Theorem 7.1. In this case, Theorem 8.1 follows by observing that for every $f\in\Lambda_{p}(\mathrm{X})$, $\min(|f|,1)\in L_{p}(\mathrm{X})$.
We conclude this section with an almost-Lebesgue differentiation theorem, being the counterpart of [11, Theorem 1.5.1] in the present theory. After a careful inspection of the proof of [11, Theorem 1.5.1], one concludes that it can be adapted to the present case since we have at our disposal the approximation of functions in $\Lambda_{p}(\mathrm{X})$ by smooth functions, and also a Chebyshev’s inequality – identity (7.6).
Theorem 8.3.
Let $f$ be a function in $\Lambda_{1}(\mathbb{R}^{d})$. Then, for almost every $x\in\mathbb{R}^{d}$ we have
$\lim_{\delta\to 0}\frac{1}{|B(x,\delta)|}\|(f-f(x))\chi_{B(x,\delta)}\|_{\alpha}=0.$ (8.3)
Remark 8.4.
If $f\in\Lambda_{1}(\mathbb{R}^{d})$, then $\min(|f|,1)\in L_{1}(\mathbb{R}^{d})$, and hence, by the Lebesgue differentiation theorem,
$\lim_{\delta\to 0}\frac{1}{|B(x,\delta)|}\|(\min(|f|,1)-\min(|f(x)|,1))\chi_{B(x,\delta)}\|_{1}=0$
which does not necessarily imply (8.3). However, if $f\in L_{1}(\mathbb{R}^{d})$, then (8.3) follows from the classical theorem since
$\|(f-f(x))\chi_{B(x,\delta)}\|_{\alpha}\leq\|(f-f(x))\chi_{B(x,\delta)}\|_{1}.$
@@ 9. Local boundedness, local convexity and duality
In the previous sections, we observed that $\Lambda_{p}(\mathbb{R}^{d})$ shares some similarities with $L_{p}(\mathbb{R}^{d})$. Since $L_{p}(\mathbb{R}^{d})$ is a normed space, it is locally bounded and locally convex. Moreover, its dual is nontrivial as it is isomorphic to $L_{q}(\mathbb{R}^{d})$ with $q=p/(p-1)$. In stark contrast to this, we will prove that the balls in $\Lambda_{p}(\mathbb{R}^{d})$ centered at 00 are neither topologically bounded nor convex, and hence $\Lambda_{p}(\mathbb{R}^{d})$ is neither locally bounded nor locally convex. Furthermore, the dual space of $\Lambda_{p}(\mathbb{R}^{d})$ is trivial.
@@
$\|f_{n}\|_{\alpha_{p}}^{p}=\int_{\mathbb{R}^{d}}|\min(n\chi_{E_{n}},1)|^{p}\,\mathrm{d}x=\int_{E_{n}}|\min(n,1)|^{p}\,\mathrm{d}x=|E_{n}|=\left(\frac{\varepsilon}{2}\right)^{p}$
@@
$\|(1/n)f_{n}\|_{\alpha_{p}}^{p}=\int_{\mathbb{R}^{d}}|\min(\chi_{E_{n}},1)|^{p}\,\mathrm{d}x=|E_{n}|=\left(\frac{\varepsilon}{2}\right)^{p}$
which does not converge to 00 as $n\to\infty$. This proves that $B_{\varepsilon}$ is not topologically bounded, which establishes the result. ∎
@@
The space $\Lambda_{p}(\mathbb{R}^{d})$ is not locally locally convex.
@@
Let $\varepsilon>0$ and $B_{\varepsilon}$ be the ball in $\Lambda_{p}(\mathbb{R}^{d})$ centered at 00 with radius $\varepsilon$. We prove that $B_{\varepsilon}$ is not convex. For a fixed $K\in\mathbb{N}$, let $g_{K}$ be given by
$g_{K}=\frac{1}{K}\sum_{n=1}^{K}f_{n}$
where $f_{n}$ is as in the proof of the previous result. We have
$\|g_{K}\|_{\alpha_{p}}^{p}=\sum_{n=1}^{K}\int_{E_{n}}|\min(n/K,1)|^{p}\,\mathrm{d}x=\left(\frac{\varepsilon}{2}\right)^{p}\frac{1}{K^{p}}\sum_{n=1}^{K}n^{p}.$
Choosing $K$ large enough so that $(1/K^{p})\sum_{n=1}^{K}n^{p}>2^{p}$ yields that $g_{K}$ does not belong $B_{\varepsilon}$. Since $g_{K}$ is a convex combination of elements in $B_{\varepsilon}$, it follows that $B_{\varepsilon}$ is not convex. Therefore, the only nontrivial open convex subset of $\Lambda_{p}(\mathbb{R}^{d})$ is $\Lambda_{p}(\mathbb{R}^{d})$ itself, and hence $\Lambda_{p}(\mathbb{R}^{d})$ is not locally convex. ∎
@@
Suppose, towards a contradiction, that there exists $\varphi\in\Lambda_{p}(\mathbb{R}^{d})^{*}$ and $f\in\Lambda_{p}(\mathbb{R}^{d})$ such that $\varphi(f)\neq 0$. By linearity, $\varphi$ is surjective, and by continuity, the set $\varphi^{-1}(-1,1)$ is open. Moreover, by linearity again, $\varphi^{-1}(-1,1)$ is convex, and so $\varphi^{-1}(-1,1)=\Lambda_{p}(\mathbb{R}^{d})$ by the previous proposition, which contradicts the surjectivity. The result follows. ∎
@@
We use the notion of $\mathrm{F}$-norm considered in [16].
@@
Let $V$ be a real vector space. An $\mathrm{F}$-norm on $V$ is a map $\|\cdot\|:V\to\mathbb{R}$ satisfying the following conditions:
@@
The map $\|\cdot\|_{\alpha_{p}}:\Lambda_{p}(\mathrm{X})\to[0,\infty)$ is an $\mathrm{F}$-norm.
@@
Acknowledgments
The author thanks João Paulos and Olena Karlova for helpful discussions. This work was supported by the Austrian Science Fund (FWF), project number 10.55776/F65.

### v3

@@ 7. Local boundedness, local convexity and duality
In the previous sections, we observed that $\Lambda_{p}(\mathbb{R}^{d})$ shares some similarities with $L_{p}(\mathbb{R}^{d})$. Since $L_{p}(\mathbb{R}^{d})$ is a normed space, it is locally bounded and locally convex. Moreover, its dual is nontrivial as it is isomorphic to $L_{q}(\mathbb{R}^{d})$ with $q=p/(p-1)$. In stark contrast to this, we prove that the balls in $\Lambda_{p}(\mathbb{R}^{d})$ centered at 00 are neither topologically bounded nor convex, and hence $\Lambda_{p}(\mathbb{R}^{d})$ is neither locally bounded nor locally convex. Furthermore, the dual space of $\Lambda_{p}(\mathbb{R}^{d})$ is trivial.
@@
$\|f_{n}\|_{\alpha_{p}}^{p}=\int_{\mathbb{R}^{d}}\min(n\chi_{E_{n}},1)^{p}\,\mathrm{d}x=\int_{E_{n}}\min(n,1)^{p}\,\mathrm{d}x=|E_{n}|=\left(\frac{\varepsilon}{2}\right)^{p}$
@@
$\|(1/n)f_{n}\|_{\alpha_{p}}^{p}=\int_{\mathbb{R}^{d}}\min(\chi_{E_{n}},1)^{p}\,\mathrm{d}x=|E_{n}|=\left(\frac{\varepsilon}{2}\right)^{p}$
which does not converge to 00 as $n\to\infty$. This proves that $B_{\varepsilon}$ is not topologically bounded, and since $\varepsilon$ is arbitrary, the result follows. ∎
@@
The space $\Lambda_{p}(\mathbb{R}^{d})$ is not locally convex.
@@
Suppose, toward a contradiction, that $\Lambda_{p}(\mathbb{R}^{d})$ is locally convex. Then, given $R>0$, there exists a convex neighborhood $C$ of zero contained in the ball $B_{R}$ of radius $R$ centered at zero. Let $20<\varepsilon<$ 2 be such that
$B_{\varepsilon}\subseteq C\subseteq B_{R}$
and let $f_{n}=n\chi_{E_{n}}$, $n\in\mathbb{N}$, be as in the proof of the previous result. Each $f_{n}$ belongs to $B_{\varepsilon}$, and therefore to $C$ as well. Let $K\in\mathbb{N}$ and define $g_{K}$ by
$g_{K}=\frac{1}{K}\sum_{n=1}^{K}f_{n}\,.$
Since $\varepsilon<2$, we have $E_{n}\cap E_{m}=\emptyset$ for $n\neq m$, and hence
$\|g_{K}\|_{\alpha_{p}}^{p}=\sum_{n=1}^{K}\int_{E_{n}}\min(n/K,1)^{p}\,\mathrm{d}x=\left(\frac{\varepsilon}{2}\right)^{p}\frac{1}{K^{p}}\sum_{n=1}^{K}n^{p}\,.$
Choosing $K$ large enough so that $(1/K^{p})\sum_{n=1}^{K}n^{p}>R^{p}$ yields that $g_{K}$ does not belong to $B_{R}$, and hence does not belong to $C$. This contradicts the convexity of $C$ given that $g_{K}$ is a convex combination of some of its elements. Consequently, $\Lambda_{p}(\mathbb{R}^{d})$ is not locally convex. ∎
@@
Assume, by contradiction, that there exists $\varphi\in\Lambda_{p}(\mathbb{R}^{d})^{*}$ and $f\in\Lambda_{p}(\mathbb{R}^{d})$ such that $\varphi(f)\neq 0$, and let $I\subsetneq\mathbb{R}$ be an open convex set containing zero. By linearity, $\varphi$ is surjective, and by continuity, the set $\varphi^{-1}(I)$ is open. Moreover, by linearity again, $\varphi^{-1}(I)$ is convex, and so $\varphi^{-1}(I)=\Lambda_{p}(\mathbb{R}^{d})$ by the previous proposition, which contradicts the surjectivity. The result follows. ∎
@@
We use the notions of $\mathrm{F}$-norm and $\mathrm{F}$-seminorm considered in [16, 11].
@@
Let $V$ be a real vector space. An $\mathrm{F}$-seminorm on $V$ is a functional $\|\cdot\|:V\to\mathbb{R}$ satisfying the following conditions:
@@
For $p\geq 1$, the functional $\|\cdot\|_{\alpha_{p}}=\|\min(|\cdot|,1)\|_{p}$ is an $\mathrm{F}$-norm on $\Lambda_{p}(\mathrm{X})$.
@@
Corollary A.3.
For $p\geq 1$ and $F\in\Sigma$ with $\mu(F)<\infty$, the functional $\|\cdot\|_{\alpha_{p},F}=\|\min(|\cdot|,1)\chi_{F}\|_{p}$ is an $\mathrm{F}$-seminorm on $L_{0}(\mathrm{X})$.